OpenStudy (anonymous):

Where is the removable discontinuity of f(x)=(x+5)/(2x^2+3x-10) located a.x=-5 b.x = –2 c.x = 2 d.x = 5

OpenStudy (michele_laino):

please can you check your formula of f(x)?

OpenStudy (michele_laino):

is the denominator, like this: \[2{x^2} + 3x - 10\]

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

the solution of this equation: \[2{x^2} + 3x - 10 = 0\] are: \[\begin{gathered} x = \frac{{ - 3 \pm \sqrt {{3^2} - 4 \times 2 \times \left( { - 10} \right)} }}{{2 \times 2}} = \hfill \\ \hfill \\ = \frac{{ - 3 \pm \sqrt {9 + 80} }}{4} = \frac{{ - 3 \pm \sqrt {89} }}{4} \hfill \\ \end{gathered} \] am I right?

OpenStudy (michele_laino):

I have applied the standard formula: \[x = \frac{{ - b \pm \sqrt {{b^2} - 4 \times a \times c} }}{{2 \times a}}\]

OpenStudy (michele_laino):

I think that the denominator can be this: \[{x^2} + 3x - 10\] is it possible?